ER004A

PI: Ruscica
Title: BL Lacs

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 15:30(15/02)-02:00(16/02)
Stations: EbHhJbMcNtOnSmShTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

JIVE (venturi at Sat Feb 15 21:29:18 EST 1997) Success
Medicina station: no known problem appeared

Onsala (John Conway at Sun Feb 16 16:27:10 EST 1997) Success
No known problesm

HartRAO (Jonathan Quick at Mon Feb 17 02:31:45 EST 1997) Minor failures
Read heads failing - Write heads seemingly still OK

Westerbork (Tony Foley at Tue Feb 18 10:04:18 EST 1997) Success
Looks OK

Torun (Roman Feiler at Wed Feb 19 17:53:57 EST 1997) Minor failures
19:00-19:30 off source, BBC7 still not operational (BBC8 used instead).

Jodrell (Alastair Gunn at Fri Feb 21 03:30:13 EST 1997) Failure
POLARISATIONS ARE CROSSED. Therefore no useful data.

Shanghai (Huang Xinyong at Wed Mar 5 04:43:56 EST 1997) Success
Clock offset = -0.09 msec. No problems, weather fine.

Calibration (Fredrik T Rantakyro at Thu May 8 11:30:48 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name er4a.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.